### v2

# CrossIn: An Efficient Instruction Tuning Approach for Cross-Lingual Knowledge Alignment

Abstract
Multilingual proficiency in large language models (LLMs) presents a significant challenge due to the uneven distribution of training data and the English-centric focus during instruction tuning. To mitigate these issues, we introduce CrossIn (Cross-lingual Instruction Tuning), which utilizes two distinct types instruction tuning datasets: the Complex Task Dataset (CTD), rich in diverse, high-quality and logical tasks like math and coding, and the Linguistic Uniformity Dataset (LUD), consisting of easier-to-translate, linguistically uniform tasks. CrossIn merges cross-lingual instruction data from CTD with machine-translated data from LUD to reinforce knowledge alignment during instruction tuning. This strategy allows us to enhance reasoning capabilities without compromising knowledge consistency across languages. We also present a multi-task benchmark to evaluate CrossIn, with results showing substantial improvements in performance across languages and tasks. This demonstrates the benefits of integrating cross-lingual data and translation in enhancing multilingual consistency and accuracy during instruction tuning.  [^1] We plan to release our datasets and models following the conclusion of the anonymity period.
CrossIn: An Efficient Instruction Tuning Approach for Cross-Lingual Knowledge Alignment
[…]
1 Introduction
The advancement of large language models (LLMs) like ChatGPT achiam2023gpt and Gemma team2023gemini has been a game-changer in the field of natural language processing (NLP), revolutionizing tasks such as language generation and commonsense reasoning naveed2024comprehensive. Nevertheless, most state-of-the-art LLMs are English-centric, and their performance on non-English languages is usually suboptimal, especially on languages that are dissimilar to English blevins2022language; mehrabi2022survey; gao2024multilingual. This challenge mainly stems from the imbalanced distribution of multilingual data at both the pre-training and instruction tuning stages. The exposure bias toward major languages results in an imbalanced capability, where models excel in languages with plentiful data while under-performing in those with limited resources dac2023okapi; feng2023pretraining. Bridging the language gap is a fundamental step to unlock the full potential of these general-purpose models and ensure that the benefits are accessible to people across the linguistic spectrum zhu2023multilingual.
Efforts to improve the multilingual capabilities of English-centric LLMs have involved continue pre-training using extensive language-specific datasets. Yet, mastering languages through additional pre-training could require vast amounts of data and significant computational resources le2022bloom. On the other hand, despite the limited proportion of non-English data at the pre-training stage, their absolute volume builds a solid knowledge base of various languages. In each iteration, LLMs are exposed to samples in several languages simultaneously, and the compressed representation encourages models to share linguistic features and generalize across different languages le2022bloom. However, this ability is not fully retained through the use of datasets that only include English in follow-up tuning steps.
In this paper, we explore instruction tuning with two distinct datasets: the Complex Task Dataset (CTD), which contains a variety of high-quality, hard-to-translate tasks like math and coding, and the Linguistic Uniformity Dataset (LUD), characterized by easier-to-translate, linguistically uniform tasks. Tuning with CTD often leads to lower consistency due to model forgetting luo2024empirical, while tuning with LUD may lead to lower reasoning capability due to its homogeneous nature.
To address these issues, we propose a method that enhances both the consistency of cross-lingual instruction tuning and task performance accuracy. Our approach leverages advanced tuning strategies that exploit the logical structures within tasks, thereby improving logical reasoning and model effectiveness across different languages. This method not only balances the complexities associated with language-specific nuances but also enhances overall model performance in multilingual environments. Our results demonstrate substantial improvements in multilingual proficiency and task accuracy, advancing the capabilities of language models.
To extensively evaluate the cross-lingual knowledge alignment qi2023crosslingual; wang2023seaeval, we establish a benchmark of three tasks (i.e., reading comprehension, commonsense question-answering, and logic reasoning). Consistency is measured by analyzing an LLM’s responses to the same question in different languages, and our benchmark encompasses multiple ability aspects and difficulty levels. Moreover, since exact match and F1 score cannot precisely evaluate system outputs in the generative setting, we unify all three tasks in a multiple-choice format for quantitative and reproducible evaluation. The experimental results demonstrate that our mixed cross-lingual tuning can significantly improve performance in all aspects (up to 40% relative gain), followed by a detailed analysis of the influence of data quantity on language consistency and knowledge accuracy.
The main contributions of our research are:
  • A Multi-faceted Benchmark. We present a multi-lingual, multi-capability benchmark for assessing the cross-lingual knowledge consistency of language models. In particular, we build a parallel multiple-choice version of the XQuAD dataset Artetxe:etal:2019 - Cross-XQuAD for machine comprehension, and combining it with commonsense QA and logic reasoning.
  • Mixed Cross-Lingual Instruction Tuning. We introduce CrossIn, a cross-lingual instruction tuning approach aimed at aligning knowledge across languages to stimulate the model’s full multilingual capability after pre-training. It offers a more reliable way of improving the model’s capability without loss of consistency.
  • CrossIn Data Insights. We conduct extensive experiments with representative LLMs on three tasks, and show the effectiveness of our proposed approach. We provide detailed analysis to study the optimal amount of cross-lingual data and the necessity of sample translation in enhancing models’ cross-lingual consistency.
2 Related Work
2.1 Multilingual Large Language Model
Multilingual Large Language Models (MLLMs) have experienced significant advancements in recent years. Recently, qin2024multilingual, as a comprehensive review, summarizes various methodologies for training MLLMs. BLOOM le2022bloom, Jais sengupta2023jais, and Sailor dou2024sailor are representative models that target improved multilingualism in the pretraining stage. For fine-tuning, ChatGLM employs a reward model trained under a multilingual setting zeng2022glm, while the x-LLM utilizes a translated version of the Alpaca dataset, combined with supervised translation data and instruction finetuning, to enhance the model’s multilingual capabilities zhu2023extrapolating.
Instruction tuning using English datasets has demonstrated the potential to extend zero-shot capabilities across multiple languages wei2022finetuned; chung2022scaling, however, they do have some limitation which is discussed in LABEL:sec:data_5_2. Previous research supports the notion that utilizing training sets composed of diverse languages can significantly enhance cross-lingual generalization muennighoff2023crosslingual; kew2023turning; shaham2024multilingual. Building on these insights, our work focuses on enhancing multilingual consistency by specifically targeting instruction finetuning. By optimizing the instruction processing mechanism, we aim to ensure better alignment across different languages during instruction tuning phase.
(a) Original XQuAD Dataset
(b) Cross-XQuAD Dataset Creation
Figure 1: An illustration of the dataset construction process of the Cross-XQuAD dataset. The original XQuAD dataset, although multilingual, is not adapted specifically to evaluate LLMs and their cross-lingual consistency.
2.2 Multilingual Evaluation Benchmark
Evaluating the multilingual capabilities of LLMs is crucial for their global applicability, as it ensures that these models can understand and generate text effectively across different languages. Benchmarks such as MMLU hendrycks2021measuring, TruthfulQA lin2021truthfulqa have been developed to access the general capability of the LLMs in English. XQuAD Artetxe:etal:2019 and MLQA lewis2019mlqa are popular extractive question-answering datasets that have been developed to evaluate the models’ multilingual performance. However, they focus on language-specific performance without considering the knowledge-sharing capabilities. Recently, Cross-MMLU and Cross-LogiQA wang2023seaeval are proposed to assess the multilingual capability of LLMs with an emphasis on cross-lingual consistency. However, the number of samples is limited which could generally lead to less stable evaluation results.
Dataset MCQs Number of Samples Supported Language Consistency Metric MLQA  lewis2019mlqa 5,500 (36$\times$) 7 - Eng, Zho, Spa, Vie, … NA XQuAD Artetxe:etal:2019 1,190 (7.9$\times$) 10 - Eng, Zho, Spa, Vie, … NA Cross-MMLU wang2023seaeval 150 (1$\times$) 7 - Eng, Zho, Spa, Vie, … Cross-LogiQA wang2023seaeval 176 (1.2$\times$) 7 - Eng, Zho, Spa, Vie, … Cross-XQuAD (ours) 1,190 (7.9$\times$) 4 - Eng, Zho, Spa, Vie
Table 1: A list of multilingual datasets. Multi-choice questions (MCQs) are more suitable for quantitative evaluation of large language models and evaluation for multilingual consistency. Traditional metrics such as the F1 score or Exact Match for extractive question answering can introduce unintended biases in evaluating large language models.
3 Cross-Lingual Consistency Benchmark
Since traditional multilingual evaluations often fail to cater specifically to LLMs or overlook the assessment of cross-lingual consistency in multilingual contexts, in this section, we present a targeted multilingual evaluation benchmark for cross-lingual knowledge alignment.
3.1 Datasets and Metrics
Even though there are multilingual evaluation datasets with parallel samples including MLQA lewis2019mlqa and XQuAD Artetxe:etal:2019, they are tailored for supervised extractive question-answering tasks and are unsuitable for less structured outputs of LLMs schuster2023semqa. Therefore, recently, two evaluation datasets have been developed for multilingual evaluation with cross-lingual consistency measures wang2023seaeval. Specifically, Cross-MMLU and Cross-LogiQA are designed to use multiple-choice questions, presenting parallel samples to assess the knowledge alignment capability of LLMs. These datasets focus on commonsense question answering and logical reasoning. However, as they are crafted by humans, the number of parallel samples they offer is relatively limited due to the high cost of human labor involved. This limitation could lead to less robust evaluation results.
Considering this, in our work, we enhance the cross-lingual consistency evaluation benchmark by introducing another task type: reading comprehension. Furthermore, we utilize existing high-quality parallel datasets to automatically generate new ones that are tailored for LLM evaluation. Table 1 summarizes the complete benchmark.
For evaluation metrics, we leverage the same concept as presented in wang2023seaeval. In addition to assessing the overall accuracy of each language, we also integrate cross-lingual consistency metrics, measured by “Consistency” and “AC3”. The consistency score is designed to determine whether the model provides consistent responses to parallel questions across different languages. A higher consistency score suggests that LLMs can apply common knowledge across languages and deliver uniform responses, regardless of correctness. Specifically, for the Cross-XQuAD dataset that spans four languages, the multilingual consistency metric is defined as
$M_{\{l_{1},l_{2},...,l_{s}\}}=\frac{\sum_{i=1}^{N}1\{a_{i}^{l_{1}}=a_{i}^{l_{2}}=...=a_{i}^{l_{s}}\}}{N}$ (1)
where $a_{i}^{l_{s}}$ is the answer for sample index $i$ from language $s$. Then, the consistency is computed as:
$\textit{Consistency}_{s}=\frac{\sum_{\{l_{1},l_{2},...,l_{s}\}\in C(s,g_{i})}M_{\{l_{1},l_{2},...,l_{s}\}}}{C^{s}_{4}}$ (2)
Similar to wang2023seaeval, we use $s=3$ as the default tolerant for consistency metrics, where the consistency between any three languages is computed. AC3 enhances the traditional accuracy metric by incorporating consistency, offering a more comprehensive evaluation. This approach is adopted because relying solely on consistency or accuracy does not yield a robust assessment.
$AC3_{s}=2\cdot\frac{\textit{Accuracy}\cdot\textit{Consistency}_{s}}{\textit{Accuracy}+\textit{Consistency}_{s}}$ (3)
By converting the datasets into MCQ (Multiple Choice Question) format, we can better quantify the model’s ability to select the correct answer from a set of options, thereby offering a clearer measure of its understanding and reasoning capabilities.
3.2 Cross-XQuAD Construction
Figure 1 indicates the process of constructing the Cross-XQuAD dataset from the original XQuAD dataset. It involves three steps, 1) English MCQ construction with distractive choices, 2) Parallel MCQ construction, and 3) Post-processing and quality check.
First, the original ground-truth answer from the XQuAD dataset can directly be used as the correction choice. As the XQuAD is for an extractive question-answer task, we extract the incorrect options from the provided context corpus as much as possible. Otherwise, the solution would be highly trivial with simple matching techniques. To achieve this, we prompt ChatGPT-3.5 to get the other three choices as shown in Figure 1(b).
Second, using the prepared English sample as a base, we prompt the generation of equivalent samples in the other languages. We discovered that direct translation without specific context can result in deviated interpretations due to polysemy, potentially leading to a biased evaluation. To counter this, we prompt the model with the English sample alongside its contextual counterpart in the target language to generate new samples. This approach has resulted in samples that are highly aligned across multiple languages.
Third, although LLMs can perform as a reasonable automated method for creating parallel samples li2023coannotating, we found that human intervention is essential to ensure higher accuracy. Consequently, each constructed sample undergoes a round of human review to confirm its integrity.
Following the above procedure, we construct the Cross-XQuAD dataset with 1,190 parallel samples in four languages which results in 4,760 samples in total. It is by far the largest multilingual evaluation dataset with cross-lingual consistency assessment capabilities.